# A stochastic optimization algorithm for revenue maximization in a service system with balking customers

### v1

@@ 2.2 Objective function
We are now ready to formally define the steady-state revenue generated per unit time $\Psi(p)$, given that the admission price is (permanently) $p$. The primary goal of this paper is to find an efficient algorithm to identify the revenue maximizing price $p^{*}$, defined as
@@
The following proposition characterizes $\Psi(p)$ in terms of the steady-state queueing performance. Its proof is deferred to Appendix A.1.
@@
Observe that (2.5) highlights the fundamental tradeoff inherent in the problem of admission pricing in a congested system: while a higher price increases the revenue earned per customer, it also lengthens the expected time between effective arrivals. As such, understanding the structure of the effective arrival process is central to our analysis. Crucially, it allows for the estimation of both the objective function and its gradient — quantities that form the foundation of the learning algorithm presented in Subsection 2.3.
We next impose strong concavity on $\Psi(\cdot)$ and a Lipschitz continuity on $\nabla\Psi(\cdot)$. These assumptions are instrumental in establishing the convergence rate of the price iterates generated by our algorithm (Theorem 6.4), and, in turn, in deriving a regret bound (Theorem 6.5).
@@ 2.3 Learning algorithm
In this subsection we define our stochastic gradient descent algorithm to learn the optimal price $p^{*}$. Let $\big\{T_{k}\big\}_{k\geq 1}$ be window sizes, which form a sequence of random variables that is defined in Section 4. For $k\geq 1$, define $\overline{T}_{0}:=0$ and $\overline{T}_{k}=T_{1}+\cdots+T_{k}$. Let iteration $k$ of the algorithm correspond to time interval $\big[\overline{T}_{k-1},\overline{T}_{k}\big)$. Then, for $k\geq 1$, in iteration $k$ of the algorithm the operator sets the admission price $p_{k-1}$. In the algorithm $p_{0}\in\mathcal{P}$ can be chosen freely, and $p_{k}$ is given by the recursion
@@
    $\widehat{\nabla\Psi}(p_{k-1})$ is an estimator of the gradient;
@@
    $\pi_{\mathcal{P}}(\cdot):\mathbb{R}\mapsto\mathcal{P}$ is the standard projection operator (i.e., $\pi(x):=\max(\underline{p},\min(\overline{p},x))$).
Our strategy for constructing and estimator $\widehat{\nabla\Psi}(p_{k-1})$ follows from Equation (2.5):
@@
where the estimators $\widehat{A_{\infty}}(p_{k-1})$ and $\widehat{\nabla A_{\infty}}(p_{k-1})$ are based on event counts and waiting time data. The specific construction of these estimators requires some additional notation and is detailed in Section 4.
@@ 3 The effective arrival and workload processes
Example A satisfies the requirements of Proposition 3.1, so that the associated queueing system is stable.
@@ 3.1 Sample path gradient analysis
Let $\zeta_{k}\sim\text{Uniform}[0,1]$ be, in the language of simulation methodology, the seed which generates the $k$-th effective interarrival time $A_{k}$, i.e., $A_{k}\overset{\mathrm{d}}{=}F^{-1}_{p}\big(\zeta_{k};\overline{W}_{k-1}\big)$. Recalling (2.1), we have
@@
We now introduce two assumptions concerning the gradient of the interarrival time with respect to the initial workload, which will be crucial to our analysis. Broadly speaking, these assumptions require that the sample paths generated by the recursive equations above exhibit uniformly bounded gradients with respect to the price. We subsequently verify that these conditions are satisfied in the setting presented in Example A.
@@
$0<1-\nabla_{w}F^{-1}p\big(\zeta;W\big)<\Xi[\zeta].$
@@
(Continued) In the case of Example A,
@@
The following lemma proves that Assumption 3.4 is met in the setting of Example A. It’s proof is given in Appendix A.3. Figure 4 visualizes this lemma.
@@
Figure 4: Visualization of Lemma 3.5. We simulate an M/G/1 + $H(p,V)$ queue with price $p=20$, and model parameters $\Lambda=20$, $\theta_{1}=0.1$, and $\theta_{2}=0.2$. For each effective arrival $i$, we record the value of the price derivative of the waiting time, denoted $\nabla_{p}\overline{W}_{i}$. The plot shows that $\nabla_{p}\overline{W}_{i}$ consistently lies in the interval $\left[-{\theta_{1}}{\theta_{2}},0\right]=[-0.5,0]$ for all $i$.
@@
For any $p\in\mathcal{P}$ we have $\nabla\operatorname{\mathbb{E}}\left[A_{\infty}(p)\right]=\operatorname{\mathbb{E}}\left[\nabla A_{\infty}(p)\right]$.
@@
We emphasize that Proposition 3.8 is crucial for the construction of an estimator for the gradient $\nabla\Psi(p)$, as needed in the recursive algorithm (2.6). Suppose one aims so to obtain an estimator for the gradient of the expectation of a random variable with respect to a given parameter — something that is often highly non-trivial to compute. Infinitesimal Perturbation Analysis (IPA) [11] provides a framework that justifies the interchange of the derivative and expectation operators. This allows us to build the estimator by essentially ‘differentiating the sample path’, thus yielding a pathwise gradient estimate under appropriate regularity conditions.
@@ 3.2 Coupling of systems with different initial workloads
The following proposition provides bounds on various quantities that capture the difference between the two coupled queues, typically expressed in terms of the distinct initial workloads $\overline{W}^{1}_{0}$ and $\overline{W}^{2}_{0}$. Let $\left\{A^{i}_{k}\right\}_{k\geq 1}$ for i = 1,2 be the effective interarrival times and let $\big\{\widetilde{A}^{i}_{k}\big\}_{k\geq 1}$ for $i=1,2$ be effective arrival times in the two queues. Also, let $N_{i}(\cdot)$, for $i=1,2$, be the counting process of the effective arrivals in queue $i$.

### v2

@@ 2.2 Objective function
The primary goal of this paper is to find an efficient algorithm to identify the revenue maximizing price $p^{*}$, defined as
@@
We are now ready to formally define the steady-state revenue generated per unit time $\Psi(p)$, given that the admission price is (permanently) $p$. The following proposition characterizes $\Psi(p)$ in terms of the steady-state queueing performance. Its proof is deferred to Appendix A.1.
@@
Observe that Proposition 2.2 highlights the fundamental trade-off inherent in the problem of admission pricing in a congested system. Indeed, while a higher price increases the revenue earned per customer,
  1. 1.
    it reduces the probability of joining, as in (2.4);
  2. 2.
    it also lengthens the expected time between effective arrivals, as in (2.5).
The remainder of the paper is devoted to the discussion and analysis of an algorithm for learning $p^{*}$ that relies solely on Equation (2.5). As such, understanding the structure of the effective arrival process is central to our analysis. In particular, it allows for the estimation of both the objective function and its gradient — quantities that form the foundation of the learning algorithm presented in Subsection 2.3.
We next assume strong concavity on $\Psi(\cdot)$ and a Lipschitz continuity on $\nabla\Psi(\cdot)$. These assumptions are instrumental in establishing the convergence rate of the price iterates generated by our algorithm (Theorem 6.4), and, in turn, in deriving a regret bound (Theorem 6.5).
@@ 2.3 Learning algorithm
In this subsection we define our SGD-based online algorithm to learn the optimal price $p^{*}$. Let $\{T_{k}\}_{k\geq 1}$ be a sequence of random variables representing window sizes, to be defined in Section 4. For $k\geq 1$, define $\overline{T}_{0}:=0$ and $\overline{T}_{k}=T_{1}+\cdots+T_{k}$. Let iteration $k$ of the algorithm correspond to time interval $\big[\overline{T}_{k-1},\overline{T}_{k}\big)$. Then, for $k\geq 1$, in iteration $k$ of the algorithm the operator sets the admission price $p_{k-1}$. In the algorithm $p_{0}\in\mathcal{P}$ can be chosen freely, and $p_{k}$ is given by the recursion
@@
    $\widehat{\nabla\Psi}(p_{k-1})$ is an estimator of the gradient at price $p_{k-1}$;
@@
    $\pi_{\mathcal{P}}(\cdot):\mathbb{R}\mapsto\mathcal{P}$ is the standard projection operator (i.e., $\pi_{\mathcal{P}}(x):=\max(\underline{p},\min(\overline{p},x))$).
Our strategy for constructing an estimator $\widehat{\nabla\Psi}(p_{k-1})$ follows from Equation (2.5):
@@
where the estimators $\widehat{A_{\infty}}(p_{k-1})$ and $\widehat{\nabla A_{\infty}}(p_{k-1})$ are based on event counts and waiting time data. The specific construction of these estimators requires additional notation and is detailed in Section 4.
@@ 3 The effective arrival and workload processes
Example A satisfies the requirements of Proposition 3.1, so that the associated queueing system is stable. $\spadesuit$
@@ 3.1 Sample path gradient analysis
Let $\zeta_{k}\sim\text{Uniform}[0,1]$ be the random seed used to generate the $k$-th effective interarrival time $A_{k}$, i.e., $A_{k}\overset{\mathrm{d}}{=}F^{-1}_{p}\big(\zeta_{k};\overline{W}_{k-1}\big)$. Recalling (2.1), we have
@@
We now introduce two assumptions concerning the sensitivity of the interarrival time to changes in the initial workload, which will be crucial to our analysis. Broadly speaking, these assumptions require that the sample paths generated by the recursive equations above exhibit uniformly bounded gradients with respect to the price. We subsequently verify that these conditions are satisfied in the setting presented in Example A.
@@
$0<1-\nabla_{w}F^{-1}_{p}\big(\zeta;W\big)<\Xi[\zeta].$
@@
(Continued) In the case of Example A, Assumption 3.3 is fulfilled. In particular,
@@
The following lemma proves that Assumption 3.4 is met in the setting of Example A. Its proof is given in Appendix A.3. Figure 4 visualizes this lemma.
@@
Figure 4: Visualization of Lemma 3.5. We simulate an M/G/1 + $H(p,V)$ queue with price $p=20$, and model parameters $\Lambda=20$, $\theta_{1}=0.1$, and $\theta_{2}=0.2$. For each effective arrival $i$, we record the value of the price derivative of the waiting time, denoted $\nabla_{p}\overline{W}_{i}$. The plot shows that $\nabla_{p}\overline{W}_{i}$ consistently lies in the interval $\left[-\frac{\theta_{1}}{\theta_{2}},0\right]=[-0.5,0]$ for all $i$.
@@
For any $p\in\mathcal{P}$ we have $\nabla_{p}\operatorname{\mathbb{E}}\left[A_{\infty}(p)\right]=\operatorname{\mathbb{E}}\left[\nabla_{p}A_{\infty}(p)\right]$.
@@
We emphasize that Proposition 3.8 is crucial for the construction of an estimator for the gradient $\nabla\Psi(p)$, as needed in the recursive algorithm (2.6). Suppose one aims to obtain an estimator for the gradient of the expectation of a random variable with respect to a given parameter — something that is often highly non-trivial to compute. Infinitesimal Perturbation Analysis (IPA) [11] constructs a pathwise derivative whose expectation coincides with the derivative of the stationary expectation under suitable regularity conditions. In our setting, these conditions are verified in Proposition 3.8. This allows us to build the estimator by essentially ‘differentiating the sample path’, thus yielding a pathwise gradient estimate.
@@ 3.2 Coupling of systems with different initial workloads
The following proposition provides bounds on various quantities that capture the difference between the two coupled queues, typically expressed in terms of the distinct initial workloads $\overline{W}^{1}_{0}$ and $\overline{W}^{2}_{0}$. Let $\left\{A^{i}_{k}\right\}_{k\geq 1}$ for i = 1,2 be the effective interarrival times and let $\big\{\widetilde{A}^{i}_{k}\big\}_{k\geq 1}$ for $i=1,2$ be effective arrival times in the two queues. Also, let $N_{i}(\cdot)$, for $i=1,2$, be the counting process of the effective arrivals in queue $i$. The proof of the proposition is given in Appendix A.4.
@@
The coupling featuring in Proposition 3.9 is illustrated by Figure 5.
4 Gradient estimator
Recalling the gradient estimator (2.8), the main objectives of this section are (i) to show how its components are computed and (ii) to establish almost sure convergence of the estimator when the sampling window increases.
In Section 2.3 we introduced the random variables $\{T_{k}\}_{k\geq 1}$ and $\{\overline{T}_{k}\}_{k\geq 1}$ corresponding to the window-sizes. In this section, we discuss their construction. Let $\left\{T^{*}_{k}\right\}_{k\geq 1}\in\mathbb{R}_{+}$ be a strictly increasing deterministic sequence, chosen by the service operator. This sequence is given as input to the algorithm and can be seen as a design choice. For $k\geq 1$, let $N_{k}$ be the random variable which denotes the number of effective arrivals during the $k$-th iteration, i.e., during the time interval $\big[\overline{T}_{k-1},\overline{T}_{k}\big)$, and let $\widetilde{N}_{k}=N_{1}+\cdots+N_{k}$. To avoid degenerate estimators, we would like at least one arrival in each sampling window. However, since the timing of the first arrival after time $T_{k}^{*}$ is not known, the duration of the sampling window $\big[\overline{T}_{k-1},\overline{T}_{k-1}+T_{k}\big)$ is random. Specifically, the window closes at the time of the first arrival after the deterministically chosen $T_{k}^{*}$:
  1. $\circ$
    $T_{k}\geq_{\text{a.s.}}T^{*}_{k}$, and
  2. $\circ$
Remark 4.1.
Note that the construction of the sampling windows allows for some flexibility. We primarily control the window duration; however, because the number of arrivals is random, there is a non-zero probability of observing zero arrivals, which would result in an unbounded arrival rate estimator. To address this, we extend the window until the first arrival occurs after the specified threshold $T_{k}^{*}$. Alternatively, one could fix the number of arrivals in each window as a deterministic increasing sequence. While this leads to a similar analysis, it makes the window duration random, which has distinct implications for the convergence and regret analysis in the following sections. We choose to fix window sizes as this seems more natural in a continuous time setting, and it is also more convenient in terms of the mathematical analysis of the algorithm performance.
Combining these statistics, we obtain the estimator $\widehat{\nabla\Psi}(p)$, defined by
$\widehat{\nabla\Psi}(p):=\frac{1}{\widehat{A_{\infty}}(p)}-p\frac{\widehat{\nabla A_{\infty}}(p)}{\widehat{A_{\infty}}(p)^{2}},$
as given in (2.8). Under the assumptions made above we have the following result regarding consistency of the gradient estimator.
Proposition 4.2.
Consider a window of length $T$ on which price $p$ is enforced. Then, under Assumption 2.1, $\widehat{\nabla\Psi}(p)\overset{{\rm a.s.}}{\longrightarrow}\nabla\Psi(p)$ as $T\rightarrow\infty$.
The proof relies on applying a law of large numbers for regenerative processes, and is deferred to Appendix A.5.
5 Convergence analysis of the learning algorithm
Proposition 5.1.
There exists constants $C_{1},C_{2}>0$ such that, for all $k\geq 1$,
$\nu_{k}$ $\leq C_{1},$ (5.3)
$|\beta_{k}|$ $\leq C_{2}\frac{\eta_{k-1}}{{T^{*}_{k-1}}^{2}}.$ (5.4)
Corollary 5.2.
5.1 Variability analysis
Therefore,
By definition,
It is clear that $A_{\infty}(p_{k-1})\geq_{\text{st}}\text{Exp}\big(\Lambda H\big(\underline{p},0\big)\big)\geq_{\text{st}}\text{Exp}\big(\Lambda\big)$, which implies that $\operatorname{\mathbb{E}}\big[A_{\infty}(p_{k-1})\big]>0$. Hence, by the continuous mapping theorem, for any $m\in{\mathbb{N}}$,
5.2 Bias analysis
In the analysis of the bias, the starting point is
[…]
$\leq\omega^{\rm(I)}_{k}+\overline{p}\,\omega^{\rm(II)}_{k}$, (5.6)
where
[…]
We aim to determine an upper bound on the rate at which $\beta_{k}$ vanishes as $k\to\infty$, which we do by analyzing the decay behavior of both $\omega_{k}^{\text{(I)}}$ and $\omega_{k}^{\text{(II)}}$. To this end we rely on the following two lemmas, the proofs of which are given in Appendix A.6.
Lemma 5.3.
As $k\to\infty$,
Lemma 5.4.
As $k\to\infty$,
Lemma 5.5.
As $k\to\infty,$
6 Regret Analysis
Recalling that $N\big(\overline{T}_{k-1},\overline{T}_{k}\big]=N\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]+1$, and that $p_{k-1}$ is $\operatorname{\mathscr{F}}_{k-1}$ measurable, we can decompose the above as
$R(L)=R_{1}(L)+R_{2}(L)\ =\sum_{k=1}^{L}R_{1,k}+\sum_{k=1}^{L}R_{2,k},$
where
6.1 Analysis for $R_{1}(L)$
The following lemma establishes an upper bound on $R_{1}(L)$. Define
Lemma 6.1.
For any $L$,
@@ 6.1 Analysis for $R_{1}(L)$
By the mean-value theorem for random variables [27], for a random variable $\xi_{k-1}$ taking values in $(0,1)$, and for $\widetilde{p}_{k-1}:=\xi_{k-1}p^{*}+\big(1-\xi_{k-1}\big)p_{k-1}$,
$R_{1,k}=T^{*}_{k}\operatorname{\mathbb{E}}\big[\Psi(p^{*})-\Psi(p_{k-1})\big]$ $=T^{*}_{k}\operatorname{\mathbb{E}}\Big[(p^{*}-p_{k-1})\operatorname{\mathbb{E}}\big[\nabla_{p}\Psi\big(\widetilde{p}_{k-1}\big)\ \big|p_{k-1}\big]\Big]$
$\leq T^{*}_{k}\operatorname{\mathbb{E}}\Big[\big|p^{*}-p_{k-1}\big|\operatorname{\mathbb{E}}\Big[\big|\nabla_{p}\Psi\big(\widetilde{p}_{k-1}\big)\big|\ \big|p_{k-1}\Big]\Big].$
With reference to Corollary 3.6,
$\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\overline{p})]}-\overline{p}\ G_{\max}\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\underline{p})]^{2}}\leq\nabla_{p}\Psi(p)$ $=\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(p)]}-p\frac{\nabla\operatorname{\mathbb{E}}\big[A_{\infty}(p)\big]}{\operatorname{\mathbb{E}}[A_{\infty}(p)]^{2}}\leq\frac{1}{\operatorname{\mathbb{E}}\big[A_{\infty}(\underline{p})\big]}.$
Therefore,
$R_{1,k}$ $\leq\max\left\{\frac{1}{\operatorname{\mathbb{E}}\big[A_{\infty}(\underline{p})\big]},\left|\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\overline{p})]}-\overline{p}\ G_{\max}\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\underline{p})]^{2}}\right|\right\}T^{*}_{k}\operatorname{\mathbb{E}}\big|p^{*}-p_{k-1}\big|=\sigma\ T^{*}_{k}\operatorname{\mathbb{E}}\big|p^{*}-p_{k-1}\big|.$
From this, the result follows. ∎
6.2 Analysis for $R_{2}(L)$
In this subsection we state and prove an upper bound on $R_{2}(L)$.
Lemma 6.2.
For any $L$,
Proof.
Fix $k\geq 1$ and let $\mathcal{F}_{k-1}$ denote the filtration up to time k1k-1. By construction, $p_{k-1}$ and $W(\overline{T}_{k-1})$ are $\mathcal{F}_{k-1}$ measurable. By definition of $\Psi(\cdot)$,
$\Psi(p_{k-1})T_{k}^{*}=p_{k-1}\,\mathbb{E}\!\left[N_{1}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T_{k}^{*}\big)\,\big|\,p_{k-1},W_{\infty}(p_{k-1})\right],$
where $N_{1}(\cdot)$ denotes the counting process of effective arrivals for an M/G/1+H(p,V)+H(p,V) queue with admission price $p_{k-1}$ and initial workload $W_{\infty}(p_{k-1})$. Similarly,
$\mathbb{E}\!\left[p_{k-1}N\big(\overline{T}_{k-1},\overline{T}_{k-1}+T_{k}^{*}\big)\,\big|\,\mathcal{F}_{k-1}\right]=p_{k-1}\,\mathbb{E}\!\left[N_{2}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T_{k}^{*}\big)\,\big|\,\mathcal{F}_{k-1}\right],$
where $N_{2}(\cdot)$ is defined as $N_{1}(\cdot)$ but now with initial workload $W(\overline{T}_{k-1})$. Therefore,
$R_{2,k}=p_{k-1}\,\bigg|\mathbb{E}\!\left[N_{1}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T_{k}^{*}\big)\,\big|\,p_{k-1},W_{\infty}(p_{k-1})\right]-\,\mathbb{E}\!\left[N_{2}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T_{k}^{*}\big)\,\big|\,\mathcal{F}_{k-1}\right]\bigg|$
Recalling the coupling construction in Proposition 3.9(c) with $\overline{W}^{1}_{0}=W_{\infty}(p_{k-1}),\overline{W}^{2}_{0}=W\big(\overline{T}_{k-1}\big)$, by applying the triangle inequality and using $p_{k-1}\leq\overline{p}$, we conclude
$R_{2,k}\leq\overline{p}\,\frac{\Lambda H(\underline{p},0)}{1-\gamma_{1}}\mathbb{E}\!\left[\Big|W_{\infty}(p_{k-1})-W(\overline{T}_{k-1})\Big|\right].$
Summing over $k=1,\dots,L$ completes the proof. ∎
6.3 Analysis for $R(L)$
Assumption 6.3.
For $k\geq 1$, define
$B_{k}=\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[\widehat{\nabla\Psi}(p_{k-1})-\nabla\Psi(p_{k-1})\big|\mathcal{F}_{k-1}\right]^{2}\right]^{\frac{1}{2}}.$
  1. (a)
    $\left(1+\frac{1}{k}\right)^{\alpha}\leq 1+\frac{K_{0}}{2}\eta_{k}$,
  2. (b)
    $B_{k}\leq\frac{K_{0}}{8}k^{-\alpha}$.
Theorem 6.4.
Under Assumption 6.3, for any $k\geq 1$,
Proof.
Deferred to Appendix A.7. ∎
An immediate consequence of Theorem 6.4, obtained using Jensen’s inequality, is that
$\operatorname{\mathbb{E}}\big[\big|p_{k-1}-p^{*}\big|\big]\leq C_{4}k^{-\alpha/2}.$
This gives us the rate of convergence of iterates $\{p_{k}\}_{k\geq 1}$ to $p^{*}$. This rate is in terms of the parameter $\alpha$ appearing in Assumption 6.3. By Equation (5.4), choosing a sequence $\big\{T^{*}_{k}\big\}_{k\geq 1}$ which grows fast tightens the rate at which the bias decays. This allows a potentially larger value of $\alpha$ to satisfy Assumption 6.3(b), and thereby to obtain a faster convergence rate to $p^{*}$. What this means is that using larger window sizes in the learning algorithm reduces the bias in the gradient estimators $\widehat{\nabla\Psi}(p)$, making them more accurate and enabling faster convergence to $p^{*}$. Since the service operator’s objective is to minimize regret by learning the optimal price as quickly as possible, a natural trade-off arises: larger window sizes reduce the number of iterations required but increase the time per iteration. We explore this trade-off through simulation experiments in Section 7.3.
Theorem 6.5.
Under Assumption 6.3, as $L\rightarrow\infty$,
$R(L)=\mathcal{O}\left(\sum_{k=1}^{L}T^{*}_{k}\ k^{-\alpha/2}\right).$
Proof.
Using Theorem 6.4 and Jensen’s inequality,
which implies, in conjunction with Lemma 6.1 that, as $L\to\infty$,
$R_{1}(L)=\mathcal{O}\left(\sum_{k=1}^{L}T^{*}_{k}\ k^{-\alpha/2}\right).$
Equations (A.6), (A.7), and (A.8), together with Lemma 6.2, lead to
as $L\to\infty$. Upon combining the above two displays,
as $L\to\infty$. ∎
7 Numerical experiments
In the preceding sections, we provided a comprehensive mathematical analysis of our SGD-based online learning algorithm. Section 7.1 empirically investigates how the service-time distribution influences the optimal price $p^{*}$. In Section 7.2, we study the algorithm’s performance through a series of illustrative examples that demonstrate its convergence behavior. Subsequently, in Section 7.3, we examine how different window size choices affect the convergence rate toward the optimal price, providing insight into their influence on algorithmic efficiency. We consider two examples of joining functions: the exponential form from Example A, and the following power-law form.
Example B.
Remark 7.1.
This section includes several examples based on Gamma-distributed service times. We emphasize that our notation $\mathrm{Gamma}\bigl(\alpha,\beta\bigr)$ denotes a Gamma distribution with shape parameter $\alpha$ and rate parameter $\beta$. Its mean is $\alpha/\beta$, and its variance is $\alpha/\beta^{2}$. We recall that ${\rm Exp}(\beta)$ denotes an exponentially distributed random variable with mean $1/\beta$ (and hence variance $1/\beta^{2}$). $\spadesuit$
7.1 Impact of service-time distribution
In this subsection, our goal is to illustrate how the service-time distribution $G(\cdot)$ impacts the optimal price $p^{*}$ for the system. To this end, Figure 6 presents three sets of plots of the objective function $\Psi(p)$, each corresponding to a different family of service-time distributions.
To the best of our knowledge, no closed-form expression for $\Psi(p)$ is available, not even for the simplest setting of Example A. Therefore, to obtain a reliable approximation of $\Psi(p)$ and thereby, the optimal price $p^{*}$, we simulate each example over a fine grid of prices for $N_{\text{eff}}=5\times 10^{5}$ effective arrivals each. We then compute the empirical average revenue per unit time $\widehat{\Psi}(p):=p/\widehat{A_{\infty}}(p)$, where $\widehat{A_{\infty}}(p)$ is the sample average of the $N_{\text{eff}}$ effective interarrival times. Finally, $p^{*}$ is approximated from the grid as the price which maximizes $\widehat{\Psi}(p)$. Throughout these experiments, all other model parameters are held fixed. In particular, we choose $\Lambda=20$ and $H(p,V)=\exp\bigl(-0.1p-0.2V\bigr)$.
  1. 1.
    $p^{*}$ increases whenever the mean service time increases.
  2. 2.
  3. 3.
    $\Psi(p)$, and in particular $\Psi(p^{*})$, increases whenever the mean service time increases.
  4. 4.
    $\Psi(p)$, and in particular $\Psi(p^{*})$, decreases whenever the variance of the service time increases.
7.2 Convergence of price iterates
Feature Example 1 Example 2 Example 3 Example 4
$\Lambda$ 2020 2020 2020 2020
$N_{\text{eff}}$ $5\times 10^{5}$ $5\times 10^{5}$ $5\times 10^{5}$ $5\times 10^{5}$
$p^{*}$ $\approx 9.3$ $\approx 29.0$ $\approx 29.5$ $\approx 16.5$
$\Psi(p^{*})$ $\approx 16.8$ $\approx 17.2$ $\approx 17.8$ $\approx 9.4$
$T^{*}_{k}$ $50\log(k+1)$ $50\log(k+1)$ $50\log(k+1)$ $50\log(k+1)$
$\eta_{k}$ $20/k^{0.75}$ $20/k^{0.75}$ $20/k^{0.75}$ $20/k^{0.75}$
Table 1: Model specifications for Examples 1–4.
Remark 7.2.
Certain choices of the service time distribution $G(\cdot)$ and the joining probability $H\big(p,V\big)$ give rise to plots for $\Psi(p)$ which plateau around $p^{*}$. In such cases, the algorithm naturally has more difficulty converging precisely to $p^{*}$. Example 4 illustrates this phenomenon. Figure 10(c) shows $\widehat{\Psi}(p)$ in a neighborhood of $p^{*}$, where the objective function is visibly flat. Consequently, as seen in Figure 10(b), the algorithm does not converge fully to $p^{*}$ even after 500 iterations, in contrast to the other examples in which convergence is much faster. However, it is important to note that, from a practical standpoint, this behavior is not problematic: the service operator’s primary goal is to maximize revenue. In that respect, the algorithm performs well, quickly moving into the vicinity of the optimal price, where the resulting revenue is essentially optimal. $\spadesuit$
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=150$ iterations, $p_{0}=20$.
Figure 7: SGD-based learning algorithm in action for Example 1.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=100$ iterations, $p_{0}=50$.
Figure 8: SGD-based learning algorithm in action for Example 2.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=100$ iterations, $p_{0}=10$.
Figure 9: SGD-based learning algorithm in action for Example 3.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=500$ iterations, $p_{0}=20$.
(c) $\widehat{\Psi}(p)$ versus $p$ in the neighbourhood of $p^{*}$.
Figure 10: SGD-based learning algorithm in action for Example 4.
7.3 Convergence time comparison
  • $\circ$
    We first consider in Figure 11(a) the choice $T^{*}_{k}=C\sqrt{k}$. In Section 7.2, our attention was on plots of the price iterates as functions of the iteration number. To better understand how quickly the iterates converge to $p^{*}$ in time, Figure 11(a) presents the price iterates against the actual time. The top panel of the figure displays the trajectories produced by the window choices $T^{*}_{k}=10\sqrt{k}$, $T^{*}_{k}=50\sqrt{k}$, and $T^{*}_{k}=100\sqrt{k}$, with the colored bands indicating the dispersion of the corresponding iterates. A wider band indicates more fluctuation, while a narrower band indicates stability. The bottom panel of Figure 11(a) shows the evolution of the error $\lvert p_{k-1}-p^{*}\rvert$ as a function of time. From these results, it is apparent that $C=10$ (blue curve) yields the fastest convergence, followed by the $C=50$ (red curve) and $C=100$ (green curve) converging the slowest.
    Figure 11(b) reveals that similar conclusions on the speed of convergence can be drawn for the choice $T^{*}_{k}=C\log(k+1)$ with $C\in\{10,50,100\}$.
    (a) Price iterates versus simulation time for $T^{*}_{k}=C\sqrt{k}$, $p_{0}=30$
  • $\circ$
    Varying window size rates. Next, we vary the rate at which the window sizes grow by studying the behavior of the price iterates over time for window functions of the form $T^{*}_{k}=C\,k^{\alpha}$, considering multiple values of the exponent $\alpha$. The corresponding plots are shown in Figure 12. Once again, similar to the observation in Figure 12, we observe that $\alpha=0.1$ (blue curve) yields the fastest convergence followed by $\alpha=0.3$ (red curve) and $\alpha=0.5$ (green curve) being the slowest.
The above two experiments highlight an inherent trade-off: increasing $C$ and $\alpha$ leads to longer time windows, which provide more data per iteration and thus yield more accurate gradient estimates. However, this comes at the cost of fewer updates within a fixed time budget. From Figures 1112, it becomes evident that smaller values of $C$ and $\alpha$ — i.e., slower growth of the window size — lead to faster convergence of the price iterates to the optimal price.
8 Concluding remarks
  • $\circ$
  • $\circ$
  • $\circ$
Appendix A Proofs
A.1 Proof of Proposition 2.2
Proof of Proposition 2.2.
Since the Markov chain $\big\{\overline{W}_{i}\big\}_{i\geq 1}$ is positive recurrent, and $A_{i+1}$ is a measurable function of $\big(\overline{W}_{i},\zeta_{i+1}\big)$, it follows that $\big(\overline{W}_{i},A_{i+1}\big)$ is an ergodic sequence. Hence, the ergodic theorem entails that
Then Equation (2.5) follows due to
This proves the stated. ∎
A.2 Proof corresponding to Example A
Then, as per the definition of $\Xi[\zeta]$ in Example A,
A.3 Proof of Lemma 3.5
Proof of Lemma 3.5.
First, we show via induction that for all $n\geq 0$,
Denoting
we can rewrite
which is an increasing linear map. As a result,
This completes the proof. ∎
A.4 Proof of Proposition 3.9
Proof of Proposition 3.9.